# TriVLA: A Triple-System-Based Unified Vision-Language-Action Model for General Robot Control

### v2

@@ 4.2 Dynamics Perception Module (System 3)
To incorporate extensive prior knowledge of dynamics into policy learning, we fine-tuned the open-source Stable Video Diffusion (SVD) model Blattmann et al. (2023a), which has 1.5 billion parameters, as the dynamics perception module for robot manipulation tasks. The fine-tuning utilized datasets including internet-sourced human manipulation data, internet robot manipulation data, and self-collected datasets. Then dynamics perception module $V_{\theta}$ is trained with diffusion objective, reconstructing the full video sequence $x_{0}=s_{0:T}$ in dataset $D$ from noised samples $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$:
@@
where $l_{emb}$ denotes the language feature from CLIP Radford et al. (2021).Then we froze the fine-tuned dynamics perception module in downstream action learning.
However, denoising a complete video sequence is computationally intensive and may cause open-loop control problems, as highlighted in Du et al. (2024). Furthermore, videos in raw pixel format frequently contain abundant irrelevant information that can hinder effective decision-making. To mitigate these challenges, we utilize the video diffusion model with a single forward pass. Our key insight is that the initial forward step, despite not producing a clear video, offers a coarse trajectory of future states and informative guidance. This observation is validated experimentally and illustrated in Figure 5. Specifically, we concatenate the current image $s_{0}$ with the final noised latent $q(x_{t^{\prime}}|x_{0})$ (typically white noise) and input this combination into the System 2. The latent features are then directly utilized. Previous work Xiang et al. (2023) emphasizes that up-sampling layers in diffusion models produce more effective feature representations. The feature at the $m^{th}$ up-sampling layer, with width $W_{m}$ and height $H_{m}$, is expressed as:
@@ 4.3 Policy Learning Module (System 1)
The predictive representations generated by the video diffusion model remain high-dimensional because they encode sequences of image features. To efficiently aggregate information across spatial, temporal, and multi-view dimensions, TriVLA compresses these representations into a fixed set of tokens. We initializes learnable tokens $Q_{[0:T,0:L]}$ with fixed length $T\times L$, performing spatial-temporal attention Blattmann et al. (2023b) on each corresponding frame, followed by feed-forward layers. Formally, this branch can be expressed as follows where $i$ is the index of frame:
@@
After the vision-language module (System 2) extracts vision-language tokens $Q_{vl}$, and the dynamics perception module (System 3) aggregates future dynamic features into predictive tokens $Q_{p}$, a diffusion policy is employed as the action head to generate the action sequence $a_{0}\in A$ conditioned on $Q_{vl}$ and $Q_{p}$. The aggregated tokens $Q_{vl}$ and $Q_{p}$ are integrated into the diffusion transformer blocks via cross-attention layers. The diffusion policy aims to reconstruct the original action $a_{0}$ from the noised action $a_{k}=\sqrt{\bar{\beta}_{k}}a_{0}+\sqrt{1-\bar{\beta}_{k}}\epsilon$, where $\epsilon$ denotes white noise and $\bar{\beta}_{k}$ is the noise coefficient at step $k$. This process can be interpreted as learning a denoiser $D_{\psi}$ to approximate the noise $\epsilon$. After the final DiT block, we apply an embodiment-specific action decoder $A_{d}$—a multi-layer perceptron (MLP)—to the final tokens to predict actions and minimize the following loss function:
@@ 5 Experiments
This section presents comprehensive experiments on both simulated and real-world robotic tasks to assess the performance of the proposed TriVLA. The simulation experiments utilize two distinct benchmarks specifically designed to systematically evaluate TriVLA’s effectiveness across diverse robot embodiments and manipulation tasks. Real-world experiments examine the model’s capabilities on a suite of tabletop manipulation tasks performed using Franka, Kinova, and Fair robots. These experiments are designed to demonstrate TriVLA’s capacity to acquire novel skills from a limited set of human demonstrations.
5.1 Simulation Setups and Baselines
@@
LIBERO Simulation Benchmark. We conduct evaluations on LIBERO Liu et al. (2024a), a simulation benchmark consisting of four distinct task suites: LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long. Each suite includes 10 diverse tasks, each accompanied by 50 human-teleoperated demonstrations, designed to assess the robot’s understanding of spatial relationships, object interactions, and task-specific goals. We follow the same preprocessed pipeline as in Kim et al. (2024): (1) removing pause intervals from trajectories, (2) standardizing image resolution to 256×256 pixels, and (3) applying a 180-degree rotation to all images.
MetaWorld Simulation Benchmark. Metaworld Yu et al. (2020) employs a Sawyer robot to perform diverse manipulation tasks and serves as a widely adopted benchmark for assessing the precision and dexterity of robotic policies. It comprises 50 tasks encompassing a diverse set of objects with varying difficulty levels Radosavovic et al. (2023). We utilize the official Oracle policy to collect 50 trajectories per task, forming our dataset.
@@
Direct Action Learning Method RT-1 100%ABC 0.533 0.222 0.094 0.038 0.013 0.90
@@
Future Prediction Related Method Uni-Pi 100%ABC 0.560 0.160 0.080 0.080 0.040 0.92
@@
3D Method RoboUniview 100%ABC 0.942 0.842 0.734 0.622 0.507 3.65
Ours TriVLA (ours) 100%ABC 0.968 0.924 0.868 0.832 0.818 4.37
Data Efficiency GR-1 10%ABC 0.672 0.371 0.198 0.108 0.069 1.41
@@
[…]
@@
Table 2: Performance on the MetaWorld benchmark by difficulty level (60 tasks total: 35 easy, 14 middle, 11 hard).
@@
Training Details. As detailed in Section 4, we employ a triple-system architecture. The System 2 vision-language module utilizes the pretrained Eagle-2 VLM to process visual inputs and language instructions on an NVIDIA H100 GPU. Subsequently, we fine-tune a video foundation model to serve as a manipulation-focused dynamics perception module (System 3). The datasets used for fine-tuning comprise 193,690 human manipulation trajectories Goyal et al. (2017), 179,074 robotic manipulation trajectories O’Neill et al. (2023), and downstream task videos including the official Calvin ABC, MetaWorld, and real-world videos. Considering the varying scales and quality of these datasets, we apply distinct sampling ratios following the method described in Octo. Fine-tuning the video model requires 2–3 days on 8 NVIDIA H100 GPUs. Finally, a generalist policy (System 1) is trained using the training datasets, requiring approximately 5–9 hours on four NVIDIA H100 GPUs.
Octo 75.1 $\pm$ 0.6% 78.9 $\pm$ 1.0% 85.7 $\pm$ 0.9% 84.6 $\pm$ 0.9% 51.1 $\pm$ 1.3%
Table 3: LIBERO benchmark experimental results. For each task suite (Spatial, Object, Goal, Long), we report the average success rate and standard error across 3 seeds with 500 episodes each. TriVLA achieves the best or competitive performance across all LIBERO benchmarks suites compared to baseline approaches. The bolded entries correspond to highest success rates.
Roll-out Details. The System 2 vision-language module employs a pre-trained Eagle-2 VLM to extract vision-language tokens, operating at 36.36 Hz on an NVIDIA H100 GPU. Prior studies rely on denoising high-precision videos, a process that is time-consuming and leads to low control frequency Black et al. (2023) or even open-loop control issues Du et al. (2024). In contrast, our method employs the dynamics perception module’s initial forward step, ensuring each observation passes through System 2 only once, requiring less than 85.9 ms per inference. Subsequently, the downstream policy generates an action chunk Chi et al. (2023) of 10 steps conditioned on vision-language and prediction tokens, further enhancing control frequency. This modification enables a substantially higher control frequency of 34–36 Hz on a consumer-level NVIDIA RTX H100 GPU.

### v3

@@ 4.2 Episodic Dynamics Perception (System 3)
To infuse extensive prior knowledge of dynamics into policy learning, we fine-tuned the 1.5B-parameter open-source Stable Video Diffusion (SVD) model Blattmann et al. (2023a) as the Episodic Dynamics Perception module for robot manipulation. This video diffusion model, trained on large-scale human and robotic manipulation datasets, encodes sequences of past states and predicts future scene trajectories, enabling episodic context accumulation. By leveraging diverse sources—including internet human manipulation data, robot datasets, and self-collected data—the module robustly models sequential environmental changes essential for effective policy learning. Then Episodic Dynamics Perception module $V_{\theta}$ is trained with diffusion objective, reconstructing the full video sequence $x_{0}=s_{0:T}$ in dataset $D$ from noised samples $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$:
@@
where $l_{emb}$ denotes the language feature from CLIP Radford et al. (2021).Then we froze the fine-tuned Episodic Dynamics Perception module in downstream action learning.
However, denoising a complete video sequence is computationally intensive and may cause open-loop control problems, as highlighted in Du et al. (2024). Furthermore, videos in raw pixel format frequently contain abundant irrelevant information that can hinder effective decision-making. To mitigate these challenges, we utilize the video diffusion model with a single forward pass. Our key insight is that the initial forward step, despite not producing a clear video, offers a coarse trajectory of future states and informative guidance. This observation is validated experimentally and illustrated in Figure 6. Specifically, we concatenate the current image $s_{0}$ with the final noised latent $q(x_{t^{\prime}}|x_{0})$ (typically white noise) and input this combination into the System 2. The latent features are then directly utilized. Previous work Xiang et al. (2023) emphasizes that up-sampling layers in diffusion models produce more effective feature representations. The feature at the $m^{th}$ up-sampling layer, with width $W_{m}$ and height $H_{m}$, can be expressed as:
@@ 4.3 Policy Learning Module (System 1)
Systems 2 and 3 together form the episodic world model, combining descriptive multimodal grounding with predictive temporal modeling. This combined representation allows System 1 (Policy Learning) to flexibly adapt its actions using both accumulated experience and anticipated future dynamics. The predictive representations generated by the video diffusion model remain high-dimensional because they encode sequences of image features. To efficiently aggregate information across spatial, temporal, and multi-view dimensions, TriVLA compresses these representations into a fixed set of tokens. We initializes learnable tokens $Q_{[0:T,0:L]}$ with fixed length $T\times L$, performing spatial-temporal attention Blattmann et al. (2023b) on each corresponding frame, followed by feed-forward layers. Formally, this branch can be expressed as follows where $i$ is the index of frame:
@@
After the Episodic Multimodal Perception module (System 2) extracts vision-language tokens $Q_{vl}$, and the Episodic Dynamics Perception module (System 3) aggregates future dynamic features into predictive tokens $Q_{p}$, a diffusion policy is employed as the action head to generate the action sequence $a_{0}\in A$ conditioned on $Q_{vl}$ and $Q_{p}$. The aggregated tokens $Q_{vl}$ and $Q_{p}$ are integrated into the diffusion transformer blocks via cross-attention layers. The diffusion policy aims to reconstruct the original action $a_{0}$ from the noised action $a_{k}=\sqrt{\bar{\beta}_{k}}a_{0}+\sqrt{1-\bar{\beta}_{k}}\epsilon$, where $\epsilon$ denotes white noise and $\bar{\beta}_{k}$ is the noise coefficient at step $k$. This process can be interpreted as learning a denoiser $D_{\psi}$ to approximate the noise $\epsilon$. After the final DiT block, we apply an embodiment-specific action decoder $A_{d}$, implemented as a multi-layer perceptron (MLP). This decoder processes the final tokens to predict actions and minimize the following loss function:
@@ 5 Experiments
Real-world Experimental Setups. We use a KINOVA GEN2 robot with a RealSense D455 depth camera mounted in an eye-to-hand configuration. In an indoor environment, we arranged various objects to encourage generalization of manipulation skills across different scenes. To further evaluate the episodic world model, we designed long-horizon, high-dynamic tasks that require the agent to accumulate, recall, and predict sequential multi-modal experiences. The RealSense D455 captured the entire scene and the robot’s state from both third-person and wrist perspectives.
Comparison Methods. Generalist robot policies have been extensively studied. We compare TriVLA with a representative subset of state-of-the-art and related methods, including RT-1 Brohan et al. (2022) (semantic features with EfficientNet and FiLM), Diffusion Policy Chi et al. (2023) (conditional diffusion visuomotor policy), Robo-Flamingo Li et al. (2023b) (LLM-visual integration via Flamingo Alayrac et al. (2022)), UniPi Du et al. (2024) (video prediction plus IK for action inference), MDT Reuss et al. (2024) (diffusion transformer), Susie Black et al. (2023) (goal image via InstructPix2Pix Brooks et al. (2023) and diffusion learning), GR-1 Wu et al. (2023) (autoregressive video and action policy), Robo-Uniview Liu et al. (2024b) (3D-aware encoder), Vidman Wen et al. (2024) (pre-trained video representation), Seer Tian et al. (2024) (end-to-end vision-action via inverse dynamics), and VPP Hu et al. (2024) (video diffusion temporal reasoning). These baselines encompass direct action learning, predictive modeling, and vision-language integration paradigms.
[…]
@@
Direct Action
Learning Method
RT-1 100%ABC 0.533 0.222 0.094 0.038 0.013 0.90
@@
3D Method RoboUniview 100%ABC 0.942 0.842 0.734 0.622 0.507 3.65
Future Prediction
Related Method
Uni-Pi 100%ABC 0.560 0.160 0.080 0.080 0.040 0.92
@@
TriVLA (ours) 100%ABC 0.968 0.924 0.868 0.832 0.818 4.37
Data
Efficiency
GR-1 10%ABC 0.672 0.371 0.198 0.108 0.069 1.41
@@
Table 2: LIBERO benchmark experimental results.
Octo 75.1 $\pm$ 0.6% 78.9 $\pm$ 1.0% 85.7 $\pm$ 0.9% 84.6 $\pm$ 0.9% 51.1 $\pm$ 1.3%
Figure 4: Qualitative case study of short-horizon tasks. Our TriVLA performs well on short-horizon tasks. In the real-world tasks, it leverages a triple-system architecture that unifies Episodic Multimodal Perception and Dynamics Perception—both crucial for generalizable policy learning.
Figure 5: Qualitative results of long-horizon tasks. Our TriVLA performs well on long-horizon tasks. In the CALVIN and real-world tasks, it leverages a triple-system architecture that unifies multiple systems for generalizable policy learning.
5.1 Experimental Results
Quantitative Results. Comparisons on the Calvin benchmark are presented in Table 1. Experimental results for Robo-Flamingo, Susie, GR-1, and 3D Diffuser Actors are extracted from their respective original publications. MDT results are obtained from the official implementation. RT-1 and UniPi results are sourced from Li et al. (2023b) and Black et al. (2023), respectively. We additionally executed the Diffusion Policy using the official open-source codebase with CLIP language conditioning. Our proposed TriVLA significantly outperforms previous state-of-the-art results. Remarkably, despite training on only 10% of the annotated Calvin ABC dataset, our method achieved an average task completion length of 3.46, outperforming related methods trained on the full dataset. Furthermore, TriVLA attained the highest performance on the MetaWorld benchmark comprising 60 tasks, as detailed in Table 3. TriVLA outperforms the strongest VPP baseline in average success rate.
Table 3: Performance on the MetaWorld.
@@ 5.1 Experimental Results
Quantitative results on the LIBERO benchmark are presented in Table 2,where each method is evaluated over 500 trials per task suite using 3 random seeds. The results demonstrate that TriVLA effectively adapts to LIBERO simulation tasks, achieving the best or competitive performance.
Qualitative Results. Figure 4 and  5 shows two qualitative examples of action sequences in both simulation and the real world. Given multiple consecutive instructions, TriVLA can comprehend intent and leverage prediction to complete long-horizon tasks. These results show that TriVLA supports generalizable policy learning by integrating Episodic Multimodal Perception and Episodic Dynamics Perception. Robots using TriVLA can understand complex sequential prompts, reason across extended event horizons, and adapt to dynamic scenarios for effective long-horizon task execution.
5.2 Ablation Study
Visualization of Episodic Dynamics Perception. We use a stable video diffusion model as the Episodic Dynamics Perception module. Its forward pass generates visual representations that capture both current scene information and predicted future dynamics over extended time horizons. Figure 6 visualizes ground-truth futures, single-step predictions, and full-sequence predictions on the Bridge benchmark Walke et al. (2023). Results show that full-step predictions are reasonable, while single-step representations effectively capture key motion cues—such as object and robot arm movement—benefiting downstream action learning. Overall, the Episodic Dynamics Perception module can model entire video sequences and predict future frames conditioned on current observations and instructions, demonstrating strong understanding of physical dynamics.
Figure 6: Visualization of Episodic Dynamics Perception. The red box indicates one-step prediction, the blue box corresponds to full-step prediction, and the green box marks the ground truth.
Effectiveness of Episodic Multimodal Perception Module. The System 2 Episodic Multimodal Perception module (EMP) of TriVLA is a pretrained Vision-Language Model (VLM) responsible for processing the robot’s visual perception and language instructions to interpret the environment and comprehend task goals. Experiments were conducted to evaluate the effectiveness of System 2. As summarized in Table 4, integration of the Episodic Multimodal Perception module improved performance from 4.06 to 4.37, with inference time increasing from 136 ms to 155 ms.
Table 4: Sub-system ablation studies on the CALVIN.
EMP EDP L-Policy Avg. Length $\uparrow$ Latency $\downarrow$ Params $\downarrow$
3.68 29.29ms 0.53B
4.06 115.19ms 1.87B
4.37 142.69ms 3.39B
These results suggest that integrating System 2 significantly enhances the accuracy of action generation.
Effectiveness of Episodic Dynamics Perception Module. The System 3 Episodic Dynamics Perception module (EDP) is a general-purpose video diffusion model fine-tuned on internet-sourced human and robot manipulation datasets. This module is designed to develop a controllable video generation model enhancing predictive capabilities within the manipulation domain. To assess its effectiveness, we perform an ablation study by integrating the Episodic Dynamics Perception module into the proposed TriVLA framework. The results, summarized in Table 4, were obtained using a single NVIDIA H100 GPU. Notably, incorporation of the Episodic Dynamics Perception results in a significant enhancement of performance.
6 Conclusion
TriVLA is the first framework to formalize an episodic world model within a unified triple-system architecture, drawing inspiration from cognitive neuroscience theories of episodic memory. By integrating multimodal grounding and rich temporal dynamics, TriVLA provides high-level reasoning and dynamic prediction, enabling robots to accumulate, recall, and predict sequential experiences. Experiments show that TriVLA operates efficiently and consistently outperforms state-of-the-art policy baselines. TriVLA significantly improves long-horizon reasoning, sample efficiency, and open-ended goal achievement. These results highlight the potential of episodic world model reasoning as a solid foundation for robust and generalizable robot control systems.
Ethics statement
TriVLA advances vision–language–action (VLA) research by introducing an episodic world model that enables embodied agents to accumulate, recall, and predict sequential experiences for robust, long-horizon decision-making. This capability lowers barriers to building more generalizable and intelligent robots, which can benefit applications in assistive robotics, manufacturing, and human–robot interaction. At the same time, these advances raise ethical considerations regarding potential misuse (e.g., autonomous systems acting beyond intended safety boundaries, reinforcement of social or cultural biases in pretrained vision–language models) and broader societal impacts (e.g., displacement of human labor or over-reliance on autonomous decision-making). We encourage further research into transparent evaluation, safety alignment, and bias mitigation, as well as careful consideration of the ethical implications when deploying such systems in real-world settings.
Reproducibility Statement
We are committed to ensuring the reproducibility of all results reported in this work. Upon publication, we will release the TriVLA codebase, pretrained checkpoints for all system components (vision–language model, video diffusion model, and policy network), as well as training and evaluation scripts. Detailed descriptions of the episodic world model architecture, triple-system integration, and flow-matching mechanisms are provided in Supplemental Material B, along with the hyperparameters and dataset preprocessing steps. We also specify all benchmarks, metrics, and experimental protocols used in both simulated and real-world evaluations. To facilitate independent verification, we will include environment setup instructions, hardware requirements, and example configurations, ensuring the seamless replication of our results.
Supplemental Material
To better understanding of this work, we offer additional details, analysis, and results as follows:
  • A    Use of Large Language Models (LLMs)
    In this section, we clarify the use of LLMs in our paper, which primarily served as an assistants for refining and polishing the paper during the writing process.
  • B    Implementation Details
    In this section, we present the implementation details of TriVLA and its inherent Episodic World Model, including the training procedure and rollout process.
  • C    Demo Video
    In this section, we present the performance of TriVLA on short-horizon and long-horizon tasks to verify the practical effect of the Episodic World Model.
  • D    Comparison Methods
    In this section, we select a representative subset of prior methods for comparison.
  • E    Details and More Results of Episodic Dynamics Perception
    In this section, we present detailed visualizations and results for Episodic Dynamics Perception in TriVLA. We employ a stable video diffusion model as the core module for Episodic Dynamics Perception and visualize the intermediate predictive representations through one-step and full step predictions.
  • F    Qualitative Analysis and Results.
    This section presents comprehensive experiments on simulated and real-world tasks to evaluate the TriVLA framework. The simulated experiments employ three benchmarks that encompass diverse robot embodiments and manipulation tasks. In parallel, real-world trials evaluate long-horizon tabletop manipulation using a Kinova Gen3 robotic arm.
  • G    Real-world Experiments
    In this section, we present a series of real-world experiments designed to rigorously evaluate the practical applicability, task generalization, and operational robustness of our TriVLA framework under realistic and unstructured environments.
Appendix A Use of Large Language Models (LLMs)
In preparing this paper, we used ChatGPT-4o (OpenAI) as a general-purpose writing assistance tool. Its role was strictly limited to checking and improving spelling, grammar, and sentence-level clarity. The LLM did not contribute to the conception of the research idea, experimental design, data analysis, interpretation of results, or the drafting of any substantive scientific content. All intellectual contributions, arguments, and conclusions presented in this paper are our own.
Appendix B Implementation details
B.1 Training Details.
@@ Appendix B Implementation details
As detailed in Section 4, we employ a unified triple-system architecture. System 2, the Episodic Multimodal Perception, utilizes the pretrained Eagle-2 VLM for processing visual and language inputs on an NVIDIA H100 GPU. System 3 fine-tunes a video foundation model for manipulation-centric Episodic Dynamics Perception using 193,690 human Goyal et al. (2017) and 179,074 robotic O’Neill et al. (2023) trajectories, supplemented by videos from CALVIN ABC, MetaWorld, and real-world tasks. To mitigate dataset discrepancies, we adopt dataset-specific sampling ratios following Octo. Video model fine-tuning requires 2–3 days on 8 NVIDIA H100 GPUs. The generalist policy is subsequently trained on task datasets, taking 5–9 hours on four H100 GPUs.
B.2 Roll-out Details.
The System 2 Episodic Multimodal Perception module employs a pretrained Eagle-2 VLM to extract vision-language tokens, operating at 36.36 Hz on an NVIDIA H100 GPU. In contrast to prior methods, which utilize computationally intensive video denoising—resulting in low control frequenciesBlack et al. (2023) or open-loop limitationsDu et al. (2024)—our approach processes each observation through System 2 only once, during the initial forward pass of the Episodic Dynamics Perception module, with inference latency below 85.9 ms. Subsequently, the downstream policy generates a 10-step action chunk Chi et al. (2023), enabling control frequencies of 34–36 Hz on a consumer-grade NVIDIA RTX H100 GPU.
Appendix C Demo Video
The attached video demonstrates the application of TriVLA, a triple-system architecture inspired by cognitive neuroscience, designed to enhance the capabilities of embodied agents in complex tasks. TriVLA integrates an episodic world model, enabling robots to accumulate, recall, and predict sequential multimodal experiences. This model provides the foundation for robust, adaptive control by simulating episodic memory processes. This showcases the generalization ability of the TriVLA framework. The tasks in the video highlight how TriVLA’s high-level reasoning and dynamic prediction enable robots to handle long-horizon manipulation and understand complex prompts, demonstrating its capability for sophisticated, adaptable decision-making.
The following is a detailed explanation of the tasks that TriVLA handles in the video.
C.1 Visualization of Episodic Dynamics Perception
This demo presents the Episodic Dynamics Perception capability of TriVLA in real-world scenarios. By utilizing a stable video diffusion model, TriVLA encodes the current scene and anticipates future dynamics over long time horizons. The visualization highlights ground-truth outcomes, single-step predictions, and full-sequence forecasts, demonstrating how the model captures essential motion cues such as object interactions and robotic arm trajectories. These results show that TriVLA can effectively model entire video sequences and predict future states based on current observations and task instructions, enabling a deeper understanding of physical dynamics for downstream decision-making and action planning.
C.2 Visualization of Action Trajectory in Simulation
This demo illustrates the action trajectory generation of TriVLA in a simulated environment. We showcase three representative examples of action sequences executed under multiple consecutive instructions, such as “Pull the handle to open the drawer,” “Grasp and lift the pink block,” “Use the switch to turn on the light bulb,” and “Store the grasped block in the sliding cabin.” TriVLA demonstrates its ability to comprehend complex instructions, infer underlying intent, and leverage predictive modeling to plan and execute long-horizon tasks. By combining high-level reasoning from vision-language models (VLMs) with dynamic predictive representations from video diffusion models (VDMs), TriVLA integrates world knowledge to enhance intent understanding and predicts future states to guide sequential decision-making. These results highlight how TriVLA effectively coordinates perception, reasoning, and prediction to accomplish complex, multi-step tasks in simulation.
C.3 Visualization of Action Trajectory in Short-horizon Real-world Tasks
Across the four short-horizon tasks—folding a pink towel, grasping an orange and placing it on a purple plate, pouring water from a green cup into a purple cup, and relocating a cup to the right side—the TriVLA consistently demonstrates its capability for precise, reliable, and context-aware manipulation. These tasks collectively highlight the model’s versatility: from handling deformable objects to executing accurate pick-and-place operations and controlling dynamic pouring actions. By integrating real-time visual perception with adaptive motor planning, TriVLA achieves robust short-horizon performance, ensuring accurate execution under diverse manipulation challenges. This suite of tasks underscores TriVLA’s efficiency in short-term control, where rapid perception-action coupling is critical for success.
Short-horizon Task: “Fold Towel” In this scenario, the TriVLA demonstrates its capacity to manipulate deformable objects by folding a pink towel with precision. This task requires careful handling of soft materials, demanding spatial reasoning beyond rigid object grasping. The policy leverages its perception to recognize the towel’s shape and orientation, planning an appropriate folding trajectory. Through this task, TriVLA highlights its short-horizon ability to interact with deformable objects in a controlled manner.
  • TriVLA achieves this by adjusting its grasp and fold strategy in real-time, ensuring the towel is folded along the intended line.
  • The success of this task emphasizes TriVLA’s adaptability in dealing with non-rigid objects, integrating visual cues into consistent action sequences.
Short-horizon Task: “Grasp Orange” In this scenario, the TriVLA showcases its skill in precise object manipulation by picking up an orange and placing it onto a purple plate. The task requires accurate localization and trajectory control, as the system must not only grasp the fruit securely but also transport it safely to the designated location. Through this task, TriVLA demonstrates its ability to carry out reliable short-horizon pick-and-place operations.
  • TriVLA accomplishes this by dynamically refining its grasp and movement trajectory based on real-time feedback.
  • The successful completion highlights TriVLA’s integration of perception and motion planning, enabling robust execution of targeted placement tasks.
Short-horizon Task: “Pouring” In this scenario, the TriVLA demonstrates its ability to perform liquid transfer by grasping a green cup and pouring water into a purple cup. This task requires stable control of orientation and precise alignment between the two containers, ensuring minimal spillage. The policy leverages its multimodal perception to model both rigid object states and the flow dynamics of the liquid. Through this task, TriVLA demonstrates competence in controlled pouring, a challenging short-horizon manipulation.
  • TriVLA achieves this by continuously monitoring the cup angle and relative position to regulate water flow.
  • The success of this task highlights TriVLA’s ability to coordinate fine-grained motor control with visual guidance for dynamic object interaction.
Short-horizon Task: “Pick Cup” In this scenario, the TriVLA performs a straightforward relocation task by picking up a cup from the table and placing it on the right side. While simple, this task demands accurate detection of the cup’s position and a reliable transfer motion without disrupting the environment. Through this task, TriVLA showcases its efficiency in executing basic, short-horizon manipulation.
  • TriVLA achieves this by generating a direct grasp-to-place trajectory, adapting its motion based on sensory feedback.
  • The success of this task demonstrates TriVLA’s ability to reliably execute fundamental pick-and-place operations with consistency.
C.4 Visualization of Action Trajectory in long-horizon Real-world Tasks
Long-horizon Task: “Beverage Preparation” In this scenario, the TriVLA demonstrates its competence in executing a complex, sequential task: picking up a red beverage can from the shelf, pouring the beverage into a yellow cup, inserting a straw into the yellow cup, and finally pushing the cup forward. Unlike short-horizon manipulations, this task requires the integration of multiple atomic actions into a coherent sequence, demanding sustained spatial reasoning, memory of intermediate states, and precise coordination across distinct phases. Through this task, TriVLA showcases its ability to plan and execute long-horizon activities where success depends on maintaining consistency across multiple dependent steps.
  • TriVLA achieves this by decomposing the task into sub-goals, dynamically adjusting its strategy based on real-time perception and the evolving state of the environment.
  • The successful completion highlights TriVLA’s ability to combine high-level planning with fine-grained control, ensuring the transition is seamless and reliable.
  • This task demonstrates TriVLA’s strength in long-horizon reasoning, where sustained action sequences and contextual understanding are essential for achieving complex goals.
TriVLA’s episodic world model enables robots to simulate memory processes, allowing them to accumulate sequential experiences and predict future actions. This capability helps robots adapt dynamically to changing environments and illustrates how embodied agents can reason about actions and experiences in a human-like way. It adopts a triple-system architecture that integrates episodic memory, high-level reasoning, and dynamic prediction. This unified design allows robots to understand multi-step tasks, solve complex manipulation problems, and make decisions grounded in both past experiences and anticipated outcomes.
In summary, TriVLA provides a robust framework for robots, offering spatial-temporal awareness, high-level reasoning, and adaptive control over long horizons. The model demonstrates exceptional generalization ability, enabling robots to perform tasks in diverse, complex environments.
Appendix D Comparison Methods